# FlowCast: Advancing Precipitation Nowcasting with Conditional Flow Matching

### v1

@@ Abstract
Radar-based precipitation nowcasting, the task of forecasting short-term precipitation fields from previous radar images, is a critical problem for flood risk management and decision-making. While deep learning has substantially advanced this field, two challenges remain fundamental: the uncertainty of atmospheric dynamics and the efficient modeling of high-dimensional data. Diffusion models have shown strong promise by producing sharp, reliable forecasts, but their iterative sampling process is computationally prohibitive for time-critical applications. We introduce FlowCast, the first model to apply Conditional Flow Matching (CFM) to precipitation nowcasting. Unlike diffusion, CFM learns a direct noise-to-data mapping, enabling rapid, high-fidelity sample generation with drastically fewer function evaluations. Our experiments demonstrate that FlowCast establishes a new state-of-the-art in predictive accuracy. A direct comparison further reveals the CFM objective is both more accurate and significantly more efficient than a diffusion objective on the same architecture, maintaining high performance with significantly fewer sampling steps. This work positions CFM as a powerful and practical alternative for high-dimensional spatiotemporal forecasting.
@@ 1 Introduction
This work investigates Conditional Flow Matching (CFM) (lipman2023flow; tong2024improving) as a powerful and more efficient alternative designed for rapid sampling. We introduce FlowCast, the first model to apply CFM to probabilistic precipitation nowcasting. We demonstrate that FlowCast resolves the tension between accuracy and efficiency, establishing a new state-of-the-art by exceeding the performance of leading diffusion models while offering a superior performance-cost trade-off.
@@
  • We introduce FlowCast, the first application of Conditional Flow Matching to precipitation nowcasting.
  • We establish a new state-of-the-art in probabilistic performance on two diverse radar datasets, the benchmark SEVIR dataset (veillette2020sevir) and the local ARSO dataset.
@@ 2 Related Work
Probabilistic Nowcasting.
To address uncertainty quantification, probabilistic models have become central to nowcasting. GANs (ravuri2021skilful) were an early approach for producing sharp forecasts, but diffusion models (ho2020denoising) are now state-of-the-art, offering stable training and high-quality samples. PreDiff (gao2024prediff) and LDCast (leinonen2023latent) are prominent latent diffusion models for ensemble forecasting. A notable hybrid is CasCast (gong2024cascast), which uses a deterministic model for large-scale patterns and a conditional diffusion model to refine stochastic details, and is currently the state-of-the-art for probabilistic precipitation nowcasting.
@@
While diffusion models deliver state-of-the-art results, their iterative sampling is computationally demanding. Continuous Normalizing Flows (CNFs) enable more efficient sampling. Recently, Conditional Flow Matching (CFM) (lipman2023flow) has emerged as a powerful technique for training CNFs, allowing for simulation-free training and rapid generation. It has shown strong performance in computer vision (tong2024improving; dao2023flow; jin2024pyramidal), but its potential for scientific forecasting remains unexplored. To our knowledge, this work is the first to apply CFM to probabilistic precipitation nowcasting.
@@ 3.2.2 FlowCast Architecture
The FlowCast model learns the vector field $v_{\theta}(Z_{t},t,Z_{\text{past}})$. For each data sample:
  1. 1.
    The past ($X_{\text{past}}$) and future ($X_{\text{future}}$) sequences are encoded into latent space ($Z_{\text{past}}$, $Z_{\text{future}}$).
  2. 2.
    A random noise sequence $Z_{P}\sim\mathcal{N}(0,\mathbf{I})$ and a time step $t\sim U[0,1]$ are sampled and an interpolated latent state is computed: $Z_{t}=(1-t)Z_{P}+tZ_{\text{future}}+\sigma\epsilon$.
  3. 3.
    The model predicts the vector field $\hat{v}=v_{\theta}(Z_{t},t,Z_{\text{past}})$.
  4. 4.
    The training loss is the MSE between predicted and target fields: $\mathcal{L}=\|\hat{v}-(Z_{\text{future}}-Z_{P})\|^{2}$.
@@
To generate an ensemble of forecasts:
  1. 1.
    The past radar sequence $X_{\text{past}}$ is encoded into its latent representation $Z_{\text{past}}$.
  2. 2.
    For each ensemble member, a distinct initial latent sequence $Z(0)$ is sampled from $\mathcal{N}(0,\mathbf{I})$.
  3. 3.
    The ordinary differential equation (ODE) $\frac{dZ(t)}{dt}=v_{\theta}(Z(t),t,Z_{\text{past}})$ is solved from $t=0$ to $t=1$ with an ODE solver.
  4. 4.
    The resulting forecast at $t=1$, $Z(1)$, is decoded back into pixel space: $\hat{X}_{\text{future}}=\mathcal{D}(Z(1))$.
The training and sampling procedures are illustrated in Figure 2.
Figure 2: FlowCast training and sampling procedures.
@@ 4.1.2 Evaluation
Following prior work (veillette2020sevir; gao2024prediff; gong2024cascast), we evaluate forecasts by converting radar fields to binary masks at given thresholds and computing the False Alarm Ratio (FAR), Critical Success Index (CSI), and Heidke Skill Score (HSS). We also use a max-pooled CSI over 16x16km regions (CSI-P16-M) to better assess predictions of localized extreme events, crucial for warning systems. We report the mean of these scores across all thresholds (”-M”), and also the CSI at the highest threshold for each dataset.
@@ 4.1.3 Training Details
We train our CFM model for 200 epochs using the AdamW optimizer with a learning rate of 5e-4 and a cosine scheduler. We set the standard deviation of the I-CFM probability path to a small constant $\sigma=0.01$ (tong2024improving). Model checkpoints are maintained using an exponential moving average of weights (ho2020denoising), with a decay factor of 0.999, and we keep the model checkpoint with the highest CSI-M evaluated on a subset of the validation set. The model is trained with 4 NVIDIA H100 for 7 days, with a global batch size of 12. Further implementation details are provided in Appendix A.1.
@@ 4.2 Comparison to the State of the Art
We evaluate FlowCast against three deterministic baselines: U-Net (veillette2020sevir), Earthformer (gao2022earthformer), and SimVPv2 (tan2025simvpv2), as well as two probabilistic baselines: PreDiff (gao2024prediff) and CasCast (gong2024cascast). All models are trained following their publicly released code, with the training budget fixed at 200 epochs. For probabilistic models, we adopt our evaluation protocol by selecting the checkpoint with the highest CSI-M on a validation subset, using exponential moving average weights to ensure comparability.
As shown in Table 3 and Figure 3, FlowCast sets a new state-of-the-art, achieving the highest overall CSI-M and HSS-M and the lowest CRPS. Its advantage is most pronounced at longer lead times, with a final CSI-219 score 256% higher than the best deterministic model. For detecting localized extreme events (CSI-P16-M), FlowCast is competitive with CasCast, which achieves a 2.8% higher score but with a 17.8% higher False Alarm Ratio (FAR-M). This suggests CasCast’s score may stem from over-prediction, while FlowCast better balances detection and precision.
@@
Model CRPS $\downarrow$ CSI-M $\uparrow$ CSI-P16-M $\uparrow$ HSS-M $\uparrow$ FAR-M $\downarrow$ CSI-M $\uparrow$ CSI-219 $\uparrow$
U-Net 0.0273 0.394 0.384 0.497 0.308 0.259 0.009
Earthformer 0.0252 0.411 0.407 0.518 0.285 0.280 0.016
SimVP 0.0249 0.423 0.424 0.532 0.298 0.280 0.012
PreDiff 0.0189 0.413 0.423 0.523 0.313 0.281 0.018
CasCast 0.0201 0.442 0.520 0.562 0.383 0.311 0.054
FlowCast (ours) 0.0182 0.460 0.506 0.580 0.325 0.324 0.057
@@
Figure 4 compares forecast sequences from FlowCast with the baselines. FlowCast produces sharp, perceptually realistic forecasts, avoiding the smoothness of deterministic models. More examples are provided in Appendix A.2.1.
Similarly, results for the ARSO dataset are shown in Table 4. FlowCast again achieves the best scores in almost all metrics, demonstrating its advantages on a different dataset with distinct local dynamics.
@@
Model CRPS $\downarrow$ CSI-M $\uparrow$ CSI-P16-M $\uparrow$ HSS-M $\uparrow$ FAR-M $\downarrow$ CSI-M $\uparrow$ CSI-219 $\uparrow$
U-Net 0.0264 0.399 0.432 0.505 0.371 0.260 0.011
Earthformer 0.0270 0.403 0.439 0.512 0.409 0.274 0.010
SimVP 0.0267 0.415 0.462 0.526 0.401 0.288 0.029
PreDiff 0.0211 0.369 0.411 0.471 0.400 0.241 0.010
CasCast 0.0253 0.373 0.511 0.483 0.488 0.277 0.057
FlowCast (ours) 0.0209 0.420 0.514 0.535 0.422 0.315 0.073
@@ 4.3.1 CFM objective against Diffusion
The results in Table 5 clearly demonstrate the superiority of the CFM objective. With just 10 steps, FlowCast (CFM) drastically outperforms the DDIM sampler, even a 100-step DDIM baseline, across key metrics like CRPS and CSI-M.
Table 5: Ablation study: CFM vs. diffusion objective. Results highlight the superior performance and efficiency of the CFM framework. All metrics are computed over a 12-step forecast, except ”Forecast @ +65 min” which only considers the last frame.
Forecast @ 12 steps Forecast @ +65 min
Model CRPS $\downarrow$ CSI-M $\uparrow$ CSI-P16-M $\uparrow$ HSS-M $\uparrow$ FAR-M $\downarrow$ CSI-M $\uparrow$ CSI-219 $\uparrow$ Time/Seq. (s)
CFM (10 steps) 0.0168 0.455 0.514 0.572 0.338 0.303 0.014 24
DDIM (10 steps) 0.0262 0.395 0.450 0.503 0.335 0.236 0.002 24
DDIM (50 steps) 0.0212 0.398 0.451 0.504 0.321 0.236 0.002 120
DDIM (100 steps) 0.0208 0.398 0.450 0.502 0.319 0.235 0.002 239
4.3.2 Inference Time Efficiency: Performance vs. Number of Function Evaluations
@@
Figure 5: Performance vs. efficiency trade-off. Forecast quality (CRPS $\downarrow$, CSI-M $\uparrow$) as a function of the Number of Function Evaluations (NFE). FlowCast (CFM) achieves near-optimal performance with only 3 to 10 steps, while the DDIM-based model requires 20 steps for CSI-M and 50 steps for CRPS, and degrades sharply at low NFE.

### v2

@@ Abstract
Radar-based precipitation nowcasting, the task of forecasting short-term precipitation fields from previous radar images, is a critical problem for flood risk management and decision-making. While deep learning has substantially advanced this field, two challenges remain fundamental: the uncertainty of atmospheric dynamics and the efficient modeling of high-dimensional data. Diffusion models have shown strong promise by producing sharp, reliable forecasts, but their iterative sampling process is computationally prohibitive for time-critical applications. We introduce FlowCast, the first end-to-end probabilistic model leveraging Conditional Flow Matching (CFM) as a direct noise-to-data generative framework for precipitation nowcasting. Unlike hybrid approaches, FlowCast learns a direct noise-to-data mapping in a compressed latent space, enabling rapid, high-fidelity sample generation. Our experiments demonstrate that FlowCast establishes a new state-of-the-art in probabilistic performance while also exceeding deterministic baselines in predictive accuracy. A direct comparison further reveals the CFM objective is both more accurate and significantly more efficient than a diffusion objective on the same architecture, maintaining high performance with significantly fewer sampling steps. This work positions CFM as a powerful and practical alternative for high-dimensional spatiotemporal forecasting.
@@ 1 Introduction
This work introduces FlowCast, a novel probabilistic nowcasting model built on Conditional Flow Matching (CFM) (lipman2023flow; tong2024improving), a powerful and efficient alternative designed for rapid sampling. While recent work has applied rectified flows for the deterministic refinement of blurry forecasts (fengperceptually), FlowCast is, to our knowledge, the first to successfully apply CFM as a full, noise-to-data generative model for this task. We demonstrate that FlowCast alleviates the tension between accuracy and efficiency, establishing a new state-of-the-art by exceeding the performance of leading diffusion models while offering a superior performance-cost trade-off.
We argue that CFM offers not only a computational advantage but also a superior inductive bias for this domain, specifically regarding the simplified transport of probability mass. Radar reflectivity distributions are highly multi-modal yet exhibit strong local temporal consistency. Standard diffusion models map Gaussian noise to this complex manifold via stochastic denoising or curved probability flow ODEs, often necessitating many sampling steps to resolve fine-grained structures without blurring modes. In contrast, CFM imposes a straight-line ODE prior on the generative process. This enforces the simplest possible mapping between the noise and data distributions. In the context of spatiotemporal forecasting, where temporal coherence is essential, this linear interpolation provides a much stronger and more stable prior than the winding paths of diffusion. We demonstrate that this geometric simplification allows FlowCast to maintain high fidelity with significantly fewer function evaluations.
@@
  • We introduce FlowCast, a novel full-probabilistic application of Conditional Flow Matching to precipitation nowcasting.
  • We establish a new state-of-the-art in both probabilistic performance and predictive accuracy on two diverse radar datasets, the benchmark SEVIR dataset (veillette2020sevir) and the local ARSO dataset.
@@ 2.1 Deterministic Nowcasting
2.2 Probabilistic Nowcasting
To address uncertainty quantification, probabilistic models have become central to nowcasting, aiming to sample from the full distribution of future states.
GANs and Diffusion.
GANs (ravuri2021skilful) were an early approach for producing sharp forecasts but suffer from training instability. Diffusion models (ho2020denoising) have recently emerged as the state-of-the-art, offering stable training and high-quality samples. PreDiff (gao2024prediff) and LDCast (leinonen2023latent) are prominent latent diffusion models for ensemble forecasting. A notable hybrid is CasCast (gong2024cascast), which uses a deterministic model for large-scale patterns and a conditional diffusion model to refine stochastic details.
@@ 2.2 Probabilistic Nowcasting
Generative modeling with flows offers an attractive alternative to diffusion. Traditional Continuous Normalizing Flows (CNFs) (chen2018neural) model data via ODEs but require expensive numerical integration during training to compute likelihoods, making them computationally prohibitive for high-dimensional spatiotemporal data. Conditional Flow Matching (CFM) (lipman2023flow) and Rectified Flows (liu2023flow) overcome this by regressing a vector field against a conditional probability path, enabling simulation-free training. However, while standard Rectified Flows typically utilize a singular conditional path (effectively $\sigma\to 0$), our application of Independent CFM (I-CFM) incorporates a Gaussian probability path with $\sigma>0$. This ”thickens” the training trajectory, providing crucial regularization that stabilizes the learning of the vector field for high-dimensional data compared to the singular paths of rectified flows.
Crucially, while diffusion models rely on stochastic denoising paths that are often curved and require many sampling steps, CFM allows for learning straight-line ODE trajectories between noise and data (tong2024improving). This geometric property enforces a direct mapping that preserves temporal coherence and allows for rapid sampling. While fengperceptually recently used a rectified flow module to strictly refine deterministic forecasts, FlowCast applies CFM as a standalone probabilistic generative model. This allows it to learn the full noise-to-data distribution and capture multimodal uncertainty without relying on a deterministic base forecast.
@@ 3.2.2 FlowCast Architecture
The FlowCast model learns the vector field $v_{\theta}(Z_{t},t,Z_{\text{past}})$. The complete training procedure is detailed in Algorithm 1.
Algorithm 1 FlowCast Training Process
0: Dataset $\mathcal{D}$, Pre-trained VAE Encoder $\mathcal{E}$, FlowCast Model $v_{\theta}$, standard deviation $\sigma$.
1:Initialize model parameters $\theta$
2:while not converged do
3:  Sample a batch of radar sequences $(X_{\text{past}},X_{\text{future}})\sim\mathcal{D}$
4:  Encode sequences into latent space: $Z_{\text{past}}\leftarrow\mathcal{E}(X_{\text{past}})$ and $Z_{\text{future}}\leftarrow\mathcal{E}(X_{\text{future}})$
5:  Sample prior noise $Z_{P}\sim\mathcal{N}(0,I)$, time $t\sim\mathcal{U}(0,1)$ and path Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$
6:  Compute interpolated latent state: $Z_{t}\leftarrow(1-t)Z_{P}+tZ_{\text{future}}+\sigma\epsilon$
7:  Compute target vector field: $u_{t}\leftarrow Z_{\text{future}}-Z_{P}$
8:  Predict vector field: $\hat{v}\leftarrow v_{\theta}(Z_{t},t,Z_{\text{past}})$
[…]
10:  Gradient step: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
11:end while
@@
To generate an ensemble of forecasts, we solve the learned ODE starting from noise, using the Euler method (Hairer1993-za). This process is described in Algorithm 2.
Algorithm 2 FlowCast Ensemble Sampling
0: Past radar sequence $X_{\text{past}}$, Trained FlowCast Model $v_{\theta}$, VAE Encoder $\mathcal{E}$ / Decoder $\mathcal{D}$.
0: Ensemble size NN, Number of ODE steps $S$.
0: Ensemble forecasts $\{\hat{X}_{\text{future}}^{(k)}\}_{k=1}^{N}$.
1: Encode past observations: $Z_{\text{past}}\leftarrow\mathcal{E}(X_{\text{past}})$
2:for $k=1$ to NN do
3:  Sample initial state: $Z(0)\sim\mathcal{N}(0,I)$
4:  Set step size: $\Delta t\leftarrow 1/S$
5:  Set current state: $Z_{curr}\leftarrow Z(0)$
6:  for $i=0$ to S1S-1 do
7:   Current time: $t\leftarrow i\cdot\Delta t$
8:   Predict vector field: $v\leftarrow v_{\theta}(Z_{curr},t,Z_{\text{past}})$
9:   Update state (Euler step): $Z_{curr}\leftarrow Z_{curr}+v\cdot\Delta t$
10:  end for
11:  Final latent forecast: $Z(1)\leftarrow Z_{curr}$
12:  Decode to pixel space: $\hat{X}_{\text{future}}^{(k)}\leftarrow\mathcal{D}(Z(1))$
13:end for
14:return Ensemble predictions $\{\hat{X}_{\text{future}}^{(k)}\}_{k=1}^{N}$
@@ 4.1.2 Evaluation
Following prior work (veillette2020sevir; gao2024prediff; gong2024cascast), we evaluate forecasts by converting radar fields to binary masks at given thresholds and computing the False Alarm Ratio (FAR), Critical Success Index (CSI), and Heidke Skill Score (HSS). For spatial validation, we compute the max-pooled CSI and Fractions Skill Score (FSS) over $16\times 16$ km neighborhoods (CSI-M-P16 and FSS-M-P16). We report the mean of these scores across all thresholds (”-M”) to evaluate general performance. Furthermore, to rigorously assess the detection of extreme weather events, we separately report the categorical metrics specifically at the highest intensity thresholds for each dataset.
@@ 4.1.3 Training Details
We train our CFM model for 200 epochs using the AdamW optimizer with a learning rate of 5e-4 and a cosine scheduler. We set the standard deviation of the I-CFM probability path to a small constant $\sigma=0.01$ (tong2024improving). We observed that the training process was notably stable; unlike diffusion objectives which often require complex loss weighting schedules, the I-CFM objective utilizes a simple regression loss that converged robustly without extensive hyperparameter tuning. Model checkpoints are maintained using an exponential moving average of weights (ho2020denoising), with a decay factor of 0.999, and we keep the model checkpoint with the highest CSI-M evaluated on a subset of the validation set. The model is trained with 4 NVIDIA H100 for 7 days, with a global batch size of 12. Further implementation details are provided in Appendix A.1.
@@ 4.2 Comparison to the State of the Art
We evaluate FlowCast against four deterministic baselines: U-Net (veillette2020sevir), Earthformer (gao2022earthformer), Earthfarseer (wu2024earthfarsser), and SimVPv2 (tan2025simvpv2), as well as two probabilistic baselines: PreDiff (gao2024prediff) and CasCast (gong2024cascast). All models are trained following their publicly released code, with the training budget fixed at 200 epochs. For probabilistic models, we adopt our evaluation protocol by selecting the checkpoint with the highest CSI-M on a validation subset, using exponential moving average weights.
General Performance (SEVIR & ARSO). As shown in Table 3, Figure 2, Table 4 and Figure 3, FlowCast establishes a new state-of-the-art across both diverse datasets. On SEVIR, it achieves the highest overall CSI-M, FSS-M-P16, and HSS-M and the lowest CRPS, demonstrating superior probabilistic calibration. On ARSO, FlowCast outperforms all baselines in all metrics besides FAR-M. Notably, on SEVIR, while the probabilistic baseline CasCast achieves the highest CSI-P16-M, it suffers from a significantly higher FAR-M compared to FlowCast (0.383 vs. 0.325). This indicates that FlowCast strikes a superior balance between detection sensitivity and precision, avoiding the tendency to over-predict precipitation coverage. Deterministic models achieve the lowest FAR-M scores by predicting blurry fields at longer lead times, missing the most extreme events (CSI-219).
Performance on Extreme Events. To evaluate the ability of the models to detect extreme events, we report performance at the highest intensity thresholds for both datasets in Table 5. FlowCast demonstrates a decisive advantage here. On SEVIR (Threshold 219), FlowCast achieves a CSI of 0.202, outperforming the best deterministic baseline (SimVP, 0.137) by over 47% and the leading probabilistic baseline (CasCast, 0.195). The trend holds for ARSO (Threshold 39 dBZ), where FlowCast achieves the highest CSI (0.183) and HSS (0.291). Crucially, FlowCast maintains this high detection skill while achieving a lower FAR than CasCast across all extreme thresholds, confirming its ability to generate sharp, intense features without resorting to excessive false alarms.
@@
Model CRPS $\downarrow$ CSI-M $\uparrow$ CSI-P16-M $\uparrow$ FSS-P16-M $\uparrow$ HSS-M $\uparrow$ FAR-M $\downarrow$ CSI-M $\uparrow$ CSI-219 $\uparrow$
U-Net 0.0273 0.394 0.384 0.661 0.497 0.308 0.259 0.009
Earthformer 0.0252 0.411 0.407 0.686 0.518 0.285 0.280 0.016
Earthfarseer 0.0256 0.389 0.393 0.636 0.486 0.289 0.247 0.001
SimVP 0.0249 0.423 0.424 0.701 0.532 0.298 0.280 0.012
PreDiff 0.0189 0.413 0.423 0.699 0.523 0.313 0.281 0.018
CasCast 0.0201 0.442 0.520 0.763 0.562 0.383 0.311 0.054
FlowCast 0.0182 0.460 0.506 0.767 0.580 0.325 0.324 0.057
@@
Model CRPS $\downarrow$ CSI-M $\uparrow$ CSI-P16-M $\uparrow$ FSS-P16-M $\uparrow$ HSS-M $\uparrow$ FAR-M $\downarrow$ CSI-M $\uparrow$ CSI-39 $\uparrow$
U-Net 0.0264 0.399 0.432 0.659 0.505 0.371 0.260 0.011
Earthformer 0.0270 0.403 0.439 0.691 0.512 0.409 0.274 0.010
Earthfarseer 0.0280 0.368 0.406 0.588 0.463 0.358 0.233 0.004
SimVP 0.0267 0.415 0.462 0.699 0.526 0.401 0.288 0.029
PreDiff 0.0211 0.369 0.411 0.614 0.471 0.400 0.241 0.010
CasCast 0.0253 0.373 0.511 0.712 0.483 0.488 0.277 0.057
FlowCast 0.0209 0.420 0.514 0.738 0.535 0.422 0.315 0.073
Figure 3: CSI-M and CSI at the 39 DBz threshold per lead time on the ARSO dataset. FlowCast shows significant improvements over probabilistic baselines for earlier lead times, and over deterministic baselines for later lead times.
Table 5: Comparison of FlowCast with baseline models on SEVIR and ARSO datasets for extreme events using categorical metrics with the highest thresholds per dataset. All metrics are computed over a 12-step forecast.
SEVIR ARSO
CSI HSS FAR CSI HSS FAR
Model 181 219 181 219 181 219 36 39 36 39 36 39
U-Net 0.205 0.122 0.314 0.193 0.366 0.508 0.209 0.145 0.318 0.226 0.474 0.505
Earthformer 0.229 0.109 0.348 0.180 0.354 0.343 0.216 0.145 0.335 0.231 0.531 0.553
Earthfarseer 0.194 0.097 0.291 0.152 0.341 0.412 0.159 0.104 0.245 0.164 0.445 0.449
SimVP 0.244 0.137 0.365 0.220 0.370 0.404 0.238 0.162 0.362 0.254 0.507 0.540
PreDiff 0.237 0.128 0.361 0.206 0.384 0.467 0.176 0.118 0.277 0.193 0.520 0.566
CasCast 0.286 0.195 0.427 0.309 0.501 0.567 0.202 0.142 0.320 0.235 0.647 0.694
FlowCast 0.301 0.202 0.443 0.317 0.425 0.482 0.254 0.183 0.388 0.291 0.547 0.589
Figure 4 qualitatively compares forecast sequences from FlowCast with the baselines on the SEVIR dataset. FlowCast produces sharp, perceptually realistic forecasts, avoiding the smoothness of deterministic models. Compared to the best-performing probabilistic baseline CasCast, we observe more realistic precipitation patterns, especially at longer lead times. More examples, including on ARSO, are provided in Appendix A.2.
@@ 4.3.1 CFM objective against Diffusion
The results in Table 6 clearly demonstrate the superiority of the CFM objective. With a single step, FlowCast (CFM) drastically outperforms the DDIM sampler, even a 100-step DDIM baseline, across key metrics like CRPS and CSI-M, whilst being almost 100 times faster.
Table 6: Ablation study: CFM vs. diffusion objective. Results highlight the superior performance and efficiency of the CFM framework. All metrics are computed over a 12-step forecast.
Model CRPS $\downarrow$ CSI-M $\uparrow$ CSI-P16-M $\uparrow$ FSS-M-P16 $\uparrow$ HSS-M $\uparrow$ FAR-M $\downarrow$ Time/Seq. (s)
CFM (1 steps) 0.0207 0.454 0.504 0.763 0.571 0.337 2.6
CFM (10 steps) 0.0168 0.455 0.514 0.764 0.572 0.338 24
DDIM (10 steps) 0.0262 0.395 0.450 0.622 0.503 0.335 24
DDIM (50 steps) 0.0212 0.398 0.451 0.635 0.504 0.321 120
DDIM (100 steps) 0.0208 0.398 0.450 0.664 0.502 0.319 239
4.3.2 Inference Efficiency: Performance vs. Number of Function Evaluations
@@
Figure 5: Performance vs. efficiency trade-off. Forecast quality (CRPS $\downarrow$, CSI-M $\uparrow$) as a function of NFE. FlowCast (CFM) achieves near-optimal performance with only 3 to 10 steps, while the DDIM-based model requires 20 steps to 50 steps, and degrades sharply at low NFE.